OpenStudy (camryn):

Solve the equation 2over5*x - 3= 13 for x

OpenStudy (lgbasallote):

\[\frac{2}{5x-3} = 13\]that right?

OpenStudy (camryn):

No 2 over 5 then x-3=13

OpenStudy (lgbasallote):

\[\frac{2}{5}(x-3) = 13?\]

OpenStudy (camryn):

Yes now I need the answer and the work

OpenStudy (lgbasallote):

im just gonna show you the solution... multiply both sides by 5 and you get \[2(x-3) = 13 \times 5\] now divide both sides by 2 and get\[x - 3 = \frac{65}{2}\] now add 3 to both sides \[x = \frac{65}{2} + 3\] last step is to simplify so can you do that?

OpenStudy (camryn):

Now solve the equation x + 1over 3 = 2 for x.

OpenStudy (lgbasallote):

multiply both sides by 3 you will get x + 1 = 6

OpenStudy (camryn):

So x would equal 5

OpenStudy (camryn):

Right

OpenStudy (camryn):

Ok now the formula for degrees celsius is : c=5 over 9 (f - 32),where f stands for degrees Fahrenheit. Part 1: solve the equation for f and show all steps. Part 2: determine how many degrees Fahrenheit 15 degrees celsius is.

OpenStudy (camryn):

Can someone help me